OpenStudy (anonymous):

You have three $1 bills, four $5 dollar bills, and two $10 bills in your wallet. You select a bill at random. Without replacing the bill, you choose a second bill at random. Find P($10 then $1) A. 5/72 B. 2/27 C. 5/81 D. 1/12

OpenStudy (anonymous):

Can you help me with 8 others plus my like 15 after that PLZZZZZZZZZZZZZZZZZZZZZZZZZ

OpenStudy (anonymous):

Really I need your help for about 25 questions. If you'll do it with me?

OpenStudy (triciaal):

P(1) = 3/9 p(5) = 4/9 and P(10) = 2/9 P($10 then $1) = 2/9 + 3/8

OpenStudy (anonymous):

After 2/9 + 3/8 then what?

OpenStudy (anonymous):

ok triciaal i think its d or a what about you?

OpenStudy (triciaal):

multiply so 2/9*3/8 = 1/12 sure

OpenStudy (anonymous):

1/12 isnt a answer though?

OpenStudy (anonymous):

do you have a calc nearby?

OpenStudy (triciaal):

yes choice d

OpenStudy (anonymous):

d was wrong

OpenStudy (anonymous):

was a kk though

OpenStudy (triciaal):